OpenStudy (anonymous):

another integral (it is just in order of the practice list, of ones I haven't done b4)

OpenStudy (anonymous):

\[\int\limits_{ }^{ }x^3\sqrt{1-x^2}~dx\]\[u=1-x^2,~~~~~~~~~-\frac{1}{2}du=x~dx~~~~~~~~~x^2=1-u\]

OpenStudy (anonymous):

\[\int\limits_{ }^{ }(1-u)\sqrt{u}~du\]

OpenStudy (anonymous):

oh, that even I can do.

OpenStudy (anonymous):

Try this one out for size: \[\int\frac{dx}{(x+1)(x^2+1)(x^3+1)}\]

OpenStudy (anonymous):

it becomes, \[\frac{2}{3}u \sqrt{u}-\frac{2}{5}u^2\sqrt{u}+C\]

OpenStudy (anonymous):

what? I am not good at integration by partial fraction yet. I learned integration by parts online with Khan and other tutorials, but this is a bit too much for me

OpenStudy (anonymous):

Ah what about \[\int x^{1/4}\ln x\,dx~~?\]

OpenStudy (anonymous):

I will think about this one....:)

OpenStudy (anonymous):

\[\frac{4}{5}x^{5/4}\ln x-\frac{4}{5}\int\limits_{ }^{ }x^{1/4}~dx\]

OpenStudy (anonymous):

like this, and then the obvious part, right?

OpenStudy (anonymous):

(The X's canceled)

OpenStudy (anonymous):

okay, I will post the answer too

OpenStudy (anonymous):

\[\frac{4}{5}x^{5/4} \ln x-\frac{16}{25}x^{5/4}+C\]

OpenStudy (anonymous):

I had a typo

OpenStudy (anonymous):

the power was 5/4 not 4/5

OpenStudy (anonymous):

@SithsAndGiggles I take ur challenge

OpenStudy (anonymous):

I did it no?

OpenStudy (anonymous):

for the first problem

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

and the sith's problem?

OpenStudy (anonymous):

Yup, by parts is the easiest way for that one.

OpenStudy (anonymous):

@TheLOL by all means :)

OpenStudy (anonymous):

you guys can measure strengths, but Zarkon will be stronger jk

OpenStudy (anonymous):

\[A/(x+1) + Bx+c/(x^2+1) + (dx^2 + e)/(x^3+1)\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Or Expand x^3+1

OpenStudy (anonymous):

hey I just turned ORANGE

OpenStudy (anonymous):

you will be chaning score slower and slower

OpenStudy (anonymous):

@fbi2015 feel free to browse these links for extra practice. Some of the problems are really easy, but others can be tricky. http://math.mit.edu/~sswatson/pdfs/qualifying_round_2013.pdf http://math.mit.edu/~sswatson/pdfs/qualifying_round_2014.pdf

OpenStudy (anonymous):

oh......yay

OpenStudy (anonymous):

I was thinking to finish with a last prob for today.

OpenStudy (anonymous):

LOL I HAVE TRIED THAT PAPER...

OpenStudy (anonymous):

best reference

OpenStudy (anonymous):

@fbi2015 dont stop